the slope of the line: 8x - 2y = 48
Answer by nyc_function(2733)

To find the slope, write your equation in the slope-intercept form y = mx + b.
8x - 2y = 48
-2y = (-8x + 48)
y = (-8x + 48)/-2
y = 4x - 24
The slope is the number next to x or 4.